OpenStudy (anonymous):

Not sure how to work this out exactly? Take a look and help if you can. Please

OpenStudy (anonymous):

\[\frac{( 4h - 1) }{ 1+h } ; h=2\]

OpenStudy (uri):

Put 2 in place of h

OpenStudy (anonymous):

4(2) -1

OpenStudy (anonymous):

then 1+2

OpenStudy (anonymous):

Is it as simple as that? I think I'm over thinking the problem making it harder than it is

OpenStudy (anonymous):

then what ever you get from the first answer put it over the second answer

OpenStudy (uri):

\[\frac{ 4(2)-1 }{ 1+2 }\]

OpenStudy (uri):

\[\frac{ 8-1 }{ 3}\]

OpenStudy (anonymous):

7/3?

OpenStudy (uri):



OpenStudy (anonymous):

Thank you all! :)